jigglypuff314 (jigglypuff314):

Solve the following Optimization problem. Must justify results. Decimal answers are acceptable. A square and/or a rectangle are formed using 50 inches of wire. The rectangle has a base to height ratio of 3 to 1. Determine the minimum total area.

jigglypuff314 (jigglypuff314):



jigglypuff314 (jigglypuff314):

P = 50 P = 4x + 8y 50 = 4x + 8y y = (50-4x)/8 = (25-2x)/4

jigglypuff314 (jigglypuff314):

A = x^2 + 3y^2 A = x^2 + 3((25/2x)/4)^2

OpenStudy (cwrw238):

looks good so far now use differentiation

jigglypuff314 (jigglypuff314):

yeah, that's what I need help with... I've tried it twice, got different answers both times, and was wrong both times xD ehhh I think I might be missing a chain rule or something can you help? dA/dx = 2x + 3(2)((25-2x)/4)(-1/2) ?

OpenStudy (cwrw238):

in your last line that should be 25-2x - looks like a typo

jigglypuff314 (jigglypuff314):

yes, sorry, (25-2x)/4

OpenStudy (cwrw238):

well you could expand it out and get a staight forward differentiation - it would be a bit messy though

OpenStudy (cwrw238):

lol - i just seen an error - thats not right!

jigglypuff314 (jigglypuff314):

tah dah xD what is it?

OpenStudy (cwrw238):

no - in my work - i did it by expanding first

jigglypuff314 (jigglypuff314):

lol

jigglypuff314 (jigglypuff314):

I mean, my ti 89 told me the answer to the question is 66.964... but I'm supposed to be doing this mostly by hand... soooo

OpenStudy (cwrw238):

i cant see anything wrong so far you've applied the chain rule to the second term and it looks correct

OpenStudy (cwrw238):

though it looks a bit messy to solve in that form

OpenStudy (cwrw238):

if you expand you get x^2 + 3/16 ( (25 - 2x)^2 = x^2 + 3/16( 625 -100x + 4x^2) agreed?

jigglypuff314 (jigglypuff314):

ah ha! I messed up distributing a negative before xD 2x + (-3(25-2x)/4) 2x + (-75 + 6x)/4 (8x + 6x - 75) / 4 (14x - 75) /4 then x = 75/14

jigglypuff314 (jigglypuff314):

*does happy dance* w00t got it, thanks for helping @cwrw238 ! <3

OpenStudy (cwrw238):

thats 5.375 right - yes i get that too

OpenStudy (cwrw238):

yw its is right - i did it a different way and got the same answer

OpenStudy (cwrw238):

after differentiating i got 3.5x = 18.75

OpenStudy (zarkon):

make sure you "justify results"

jigglypuff314 (jigglypuff314):

yeah, doing that now @Zarkon x A --------------- 0 | 117.1875 75/14 | 66.964 root 50 | 72.105 therefore 75/14 is the minimum area

OpenStudy (zarkon):

why root 50?

jigglypuff314 (jigglypuff314):

I made x be the side of the square and the maximum the side of the square can be when the ... wait... oh when the perimeter is 50 soooo 50/4

OpenStudy (zarkon):

yes...50/4

OpenStudy (cwrw238):

or to show its a minimum you can find the second derivative and find its sign when x = 75/14

jigglypuff314 (jigglypuff314):

yea... that was the other option... but with the way I was deriving that stuff... it'll get too ugly I think xD

OpenStudy (cwrw238):

yes with my way i got dA/dx = (7/4)x^2 - 18.75 so d^A/dx^2 = (14/4)x = 3.5 x so x = (75/14) gives a positive so minimum

jigglypuff314 (jigglypuff314):

mhmm... *nods blindly* I'll take your word for it xD I don't think it would be too bad now that I look at it... but I think I prefer plugging and chugging :) thanks for you help! :D <3